# ProGeny I: a new simple stellar population generator and impact of isochrones / stellar atmospheres / initial mass functions

### v2

@@ 2.4 Initial Mass Function
The last critical part of generating an SSP (after choosing an isochrone, combining stellar atmosphere spectral libraries, and creating a scheme to interpolate them) is to weight different stellar populations with a target initial mass function (IMF). This function simply encodes the original distribution of stellar masses that were generated by a single burst of star formation, and effectively this becomes an additional weighting term when we combine different stars in an isochrone.
IMFs in general share the characteristic that you expect to produce more (by number) stars of lower mass than higher mass. Many studies have established that there is some sort of natural lower bound to the star mass allowed, but debate continues over the exact shape of the IMF and the nature of the upper mass limit (or if there is one at all, see Weidner & Kroupa, 2004). A further point of contention is how truly Universal and/or evolving the IMF might be. Is it self-similar over all time and space? Does it naturally evolve over time but not depend on the spatial environment? Does it only vary based on the spatial environment? All these questions are still regularly discussed in the modern astronomy literature (see Kroupa, 2001; Baldry & Glazebrook, 2003; van Dokkum, 2008; Meurer et al., 2009; Bastian et al., 2010; Li et al., 2023), and access to highly flexible SSPs is a vital component in better answering them.
@@ 2.5 Output Stellar Populations Library Format
Loading the ProGeny package, loading all the relevant MIST isochrones and default moderate resolution stellar atmosphere data, running the above steps via their respective ProGeny functions, and generating a new in-memory SSP that can be directly loaded into ProSpect takes roughly 3 minutes per metallicity (logZ) output per core. The creation of SSPs in ProGeny is a natively parallel process, and by default uses 8 cores. For a particular isochrone if you have as many unique isochrone metallicity values as cores available, this should mean generating a new SSP should only ever take a couple of minutes from scratch, and much less time if you are only (e.g.) updating the IMF and not re-computing the interpolation grid weights etc.
@@ 3 The ProGeny Stellar Population Library
Below we discuss the significant decisions made regarding which isochrones (3.1), stellar atmospheres (3.2), and IMFs (3.3) are included with the initial public release of ProGeny.
@@ 3.1 Isochrones
ProGeny initially comes with three sets of modern isochrones: the MIST isochrones produced with the MESA software (Dotter, 2016; Paxton et al., 2011); the PARSEC isochrones that combine Padova and Trieste standard tracks with COLIBRI evolved tracks (Bressan et al., 2012; Marigo et al., 2017); and BaSTI (Hidalgo et al., 2018). Table 2 presents the main differences in terms of coverage and sampling between the isochrones included. MIST has the largest range of ages and metallicities, but PARSEC offers the densest sampling in metallicity space. All three are discussed briefly below, but for reasons we will outline the MIST isochrones are generally preferred for creating new SSPs.
@@ 3.1.1 MIST
The Mist isochrones were obtained from the main web hosting server444waps.cfa.harvard.edu/MIST/model_grids.html. We use the pre-packaged isochrone with v / vcrit = 0.4, and downloaded the basic version of the catalogue (25 columns). The MIST version number was 1.2, and the MESA revision number was 7503.
The main advantage of the MIST isochrones compared to older varieties is that they cover all phases of stellar evolution including remnant stars, which can be important for older stellar populations since late phase evolution and remnant stars are what provide the UV upturn component of galaxy SEDs. Missing these entirely (as many older isochrones do) can cause fitting issues for older stellar populations in highly quenched systems, and previous efforts to generate SSPs had to develop their own bespoke methods to paste these remnant populations in realistically. Notably some older SSPs do not do this at all, e.g. Maraston (2005) does not include remnant populations and has no UV upturn present at all in older stellar populations (as seen in Figure 11 discussed in detail later in this work).
@@ 3.1.2 PARSEC
PARSEC is a popular modern set of isochrones that really combine the efforts of two main groups: the Padova and Trieste isochrones that cover the majority of stellar evolution (e.g. the main sequence of stellar evolution) and later phases via the COLIBRI tracks (see Bressan et al., 2012; Marigo et al., 2017, for details). The latter covers the major asymptotic giant branch (AGB) phases that are extremely important for properly explaining the luminous output of infrared light in moderately old (few Gyr) old stars. Our compilation of the PARSEC isochrones was generated interactively via the main web server 555stev.oapd.inaf.it/cmd, and combines the v1.2S version of the Padova tracks with the S37 + S35 + PR16 combination of CALIBRI tracks (see the website for details on what these ingredients mean for the final isochrones). These isochrones cover ages in 0.1 dex bins between $10^{6.6}$$10^{10.1}$ years, and metallicity in 0.1 dex bins between $10^{-2.2}$$10^{0.5}$ $\text{Z}_{\odot}$. This means they generally cover a more limited range of ages and metallicity than MIST, and with coarser binning in age but finer in metallicity. In general for SED modelling, SSPs with finer and broader age sampling is advantageous.
The main limitation with PARSEC, and why they are not the recommended default set of isochrones, is that they lack proper remnant stellar evolution tracking. This is noted on the website, with a suggestion that this will be added longer term. As and when remnant stars are added to PARSEC, we will updated the standard isochrones available in ProGeny. We note that it is possible for users to adapt the provided isochrones and bolt on their own prescriptions for remnant star evolution (e.g. the Geneva isochrones, or the appropriate part of MIST).
@@ 3.1.3 BaSTI
The Bag of Stellar Tracks and Isochrones project (BaSTI) is an alternative isochrone library that monitors the evolutionary properties of stars up until their main sequence turn off (approximately). We downloaded pre-computed isochrones from the main INAF database666basti-iac.oa-abruzzo.inaf.it/isocs.html. The various files were converted into the appropriate format for ProGeny, with a particular note that the effective surface gravity properties of the stars had to be derived from the Stefan-Boltzmann Law (which means they are assumed to be non-rotating). These isochrones cover ages in 0.1 dex bins between $10^{7.3}$$10^{10.3}$ years, and metallicity in $\sim 0.2$ dex bins between $10^{-3.2}$$10^{0.45}$ $\text{Z}_{\odot}$. This means they generally cover a more limited range of ages and metallicity than MIST and PARSEC, and with coarser binning in age but finer in metallicity. The younger age limit is particular high (at roughly 20 Myr), making BaSTI a poor choice when computing properties of extreme star bursts (where sub 10 Myr sampling is critical).
@@ 3.1.4 Comparison
It is instructive to investigate some of the major differences between the MIST, PARSEC, and BaSTI isochrones included in the first released version of ProGeny. Figure 1 is a high-level comparison of the three families of isochrones at 100 Myr, 1 Gyr and 10 Gyr (the origin of the region polygons is made clear in Appendix B Figure 13). It is immediately apparent that the PARSEC and BaSTI isochrones are missing post-AGB stellar evolution including remnants. The main sequence region (labelled ‘MS’) is very similar between the three isochrone families, but the treatment of the RGB and TP-AGB region (bottom-left) does vary quite significantly, with PARSEC capturing the pulsations is relatively more detail. The more detailed tracking of the TP-AGB with PARSEC is not surprising, since it utilises the dedicated COLIBRI libraries for detailed modelling in this regime.
@@
The impact of variation in the TP-AGB region seems to be very modest in terms of the eventual impact on the final SSPs, but the lack of remnant stars in PARSEC and BaSTI has a very notable impact on the ultraviolet emission of older (few Gyr) stellar populations. If we apply our default interpolation of stellar atmospheres onto the isochrones we can compare the spectral flux prediction for a target Chabrier IMF. The result of this is shown in Figure 2, where there is a vastly different ultraviolet prediction due to the inclusion (MIST) or lack of inclusion (PARSEC, BaSTI) of hot remnant stars.
@@
As discussed later in Section 4, there is still a huge deal of uncertainty in the details of the ultraviolet upturn for old stellar populations and exactly what processes produce it (e.g. helium burning from Wolf-Rayet stars, post-AGB stars, horizontal branch stars, and binary interactions). There is also a scarcity of empirical data informing our knowledge of ultraviolet stellar atmospheres. The main conclusion to make here is that merely the inclusion or not of hot remnant stars alone will have a dramatic effect in the ultraviolet, but almost no impact at all on the optical to infrared regimes (i.e. any data long-wards of 2000 Å).
@@
Figure 2: Comparison of 10 Gyr ultraviolet upturn predictions for MIST, PARSEC, and BaSTI isochrones. The stellar atmospheres and interpolation used are otherwise identical, so the discrepancy in the ultraviolet is purely due to the inclusion (MIST) or lack of inclusion (PARSEC / BaSTI) of hot remnant stars. Note that the PARSEC ad BaSTI spectra lie almost on top of each other in this plot.
3.2 Stellar Atmospheres
@@
Equally as important as the isochrones, stellar atmospheres are the means by which we can convert our predictions of stellar evolution models to spectra that we can then use to model observed data, either directly as spectra, or passed through filters to create broadband photometry. Unlike isochrones, which are necessarily theoretical in nature (though informed by observations), stellar atmospheres can in principle be empirically derived using observations of many stars covering a large range of intrinsic properties, or generated theoretically. The main properties we usually wish to grid over are (in rough order of significance on the resulting spectrum): effective temperature; effective surface gravity (usually referred to as logG); and metallicity (Z). More exotically, extensions are sometimes made to cover variations in stellar rotation and/or $\alpha$ enhancement (or other abundance measures).
@@
In the perfect library of stellar atmospheres, there would exist a predicted or measured spectrum for every star for every combination of properties predicted by our stellar evolution isochrones, but this is unfeasible in practice — we would need millions of such spectra. The pragmatic solution is to map out a dense grid of these properties, with higher density where stars are most commonly found on the main sequence (since that will ultimately produce more accurate composite spectra). Rarer regions of temperate, logG and Z can be covered more coarsely without biasing any resulting spectrum too much. Once we have established this dense grid, we can use the interpolation scheme (discussed above) to heuristically produce an approximate spectrum for any target combination of properties when moving along a given isochrone.
@@
Stellar population libraries exist that choose to place more emphasis on either empirical or theoretical stellar atmospheres, but the main tradeoffs are accuracy (assuming we can estimate stellar properties, observations should always be closer to ‘reality’) and coverage. In the case of coverage, observed databases of stellar atmospheres tend to be much smaller (usually numbering hundreds), and only map out a limited range of parameter values. This makes interpolating or extrapolating these libraries onto target parameters quite complex, with spline interpolation and/or kriging schemes being necessary to achieve reasonable performance. In practice nearly all observationally focussed SPLs need to supplement the empirical stellar atmospheres with extreme/rare theoretical templates (e.g. for hot stars). The counter is also true, since some observed stellar atmospheres (e.g. TP_AGB stars) cannot be accurately reproduced theoretically.
@@
Table 3: Parameter ranges of the various stellar atmosphere libraries initially includes with ProGeny. N is the total number of unique spectra.
Library Type Teff (K) logG logZ N Reference
C3K (Conroy) Base 2,000 – 50,000 -1 – 5.5 -2.1 – 0.5 8,602 Conroy et al. (2018)
PHOENIX (Husser) Base 2,300 – 12,000 0 – 6 -4 – 1 7,559 Husser et al. (2013)
PHOENIX (Allard) Extend 2,000 – 70,000 0 – 5.5 -4 – 0.5 12,045 Allard et al. (2012)
MILES (Vazdekis) Alternative 2,000 – 50,000 -1 – 5.5 -1.4 – 0.2 3,915 Vazdekis et al. (2010)
BaSeL (WLBC) Alternative 2,000 – 50,000 -1 – 5.5 -2 – 0.5 4,649 Westera et al. (2002)
ATLAS9 (Castelli) Alternative 2,000 – 50,000 -1 – 5.5 -1 – 0.5 2,735 Castelli & Kurucz (2003)
AGB (Lacon) AGB 2,000 – 4,000 NA NA 14 Lançon & Mouhcine (2002)
TMAP (Werner) Hot (inc. white dwarfs) 20,000 – 190,000 5 – 9 NA 124 Werner et al. (2003)
@@
Below we discuss the main decisions made regarding the fiducial stellar atmospheres used within ProGeny (see Table 3). It is always possible for users to supplement these with their own preferences and use the ProGeny software to generate new SSPs, but when we discuss the base ProGeny SSPs they have been constructed with a mixture of the below libraries. All stellar atmosphere spectra provided as part of the fiducial ProGeny library encode the wavelength in angstroms and the luminosity in $\delta\lambda$ units such that the integral over all wavelengths of $f(\lambda).\delta\lambda=1$. This means all spectra can be scaled directly by the predicted stellar luminosity in an isochrone to achieve the correct output spectrum. Any additional stellar atmosphere libraries added to ProGeny by users should obey the same requirements. The spectra are also re-binned to match the exact wavelength sampling of CB19, since via inspection this appeared to offer a good compromise of moderately high resolution (especially around important feature), coverage and data volume. We note that when loading stellar atmospheres it is possible to target any desired wavelength grid, with flux conserving spectral resampling handled within ProGeny(written using efficient C++code).

### v3

@@ 2.4 Initial Mass Function
The last critical part of generating an SSP (after choosing an isochrone, combining stellar spectra libraries, and creating a scheme to interpolate them) is to weight different stellar populations with a target initial mass function (IMF). This function simply encodes the original distribution of stellar masses that were generated by a single burst of star formation, and effectively this becomes an additional weighting term when we combine different stars in an isochrone (for a detailed discussion on the physics and measurements of the IMF see Kroupa et al., 2024).
IMFs in general share the characteristic that you expect to produce more (by number) stars of lower mass than higher mass. Many studies have established that there is some sort of natural lower bound to the star mass allowed, but debate continues over the exact shape of the IMF and the nature of the upper mass limit (or if there is one at all, see Weidner & Kroupa, 2004). A further point of contention is how truly Universal and/or evolving the IMF might be. Is it self-similar over all time and space? Does it naturally evolve over time but not depend on the spatial environment? Does it only vary based on the spatial environment? All these questions are still regularly discussed in the modern astronomy literature (see Kroupa, 2001; Baldry & Glazebrook, 2003; van Dokkum, 2008; Meurer et al., 2009; Bastian et al., 2010; Li et al., 2023; Haslbauer et al., 2024), and access to highly flexible SSPs is a vital component in better answering them.
@@ 2.5 Output Stellar Populations Library Format
Loading the ProGeny package, loading all the relevant MIST isochrones and default moderate resolution stellar spectrum data, running the above steps via their respective ProGeny functions, and generating a new in-memory SSP that can be directly loaded into ProSpect takes roughly 3 minutes per metallicity output per core (this timing is based on using an Apple M3 Max processor). The creation of SSPs in ProGeny is a natively parallel process, and by default uses 8 cores. For a particular isochrone if you have as many unique isochrone metallicity values as cores available, this should mean generating a new SSP should only ever take a couple of minutes from scratch, and much less time if you are only (e.g.) updating the IMF and not re-computing the interpolation grid weights etc.
The in-memory SSP can be saved directly to a single FITS file using the Rfits package (see Robotham et al., 2024). The high-level ProSpect package has a function available (speclib_FITSload) to read the FITS SSP file back into memory with the correct structure. The details of the on-disk structure are described in the ProSpect manual for interested readers, so we will not describe it here.
@@ 3 The ProGeny Stellar Population Library
Below we discuss the significant decisions made regarding which isochrones (3.1), stellar spectra (3.2), and IMFs (3.3) are included with the initial public release of ProGeny. We also outline the significant data formats for both input and output data, and how the input ingredients are computationally combined to create target SSPs.
@@ 3.1 Isochrones
ProGeny initially comes with three sets of modern isochrones: the MIST isochrones produced with the MESA software (Dotter, 2016; Paxton et al., 2011); the PARSEC isochrones that combine Padova and Trieste standard tracks with COLIBRI evolved tracks (Bressan et al., 2012; Marigo et al., 2017); and BaSTI (Hidalgo et al., 2018). Table 2 presents the main differences in terms of coverage and sampling between the isochrones included. We note that the youngest stellar population available in any current ProGeny isochrone is 100,000 yr (MIST). This is appropriate for a large range of galaxy evolution use-cases, but might not be appropriate for very extreme bursts where even younger stellar populations can dominate the spectrum.
MIST has the largest range of ages and metallicities, but PARSEC offers the densest sampling in metallicity space. All three are discussed briefly below, but for reasons we will outline the MIST isochrones are generally preferred for creating new SSPs.
@@ 3.1.1 MIST
The MIST isochrones were obtained from the main web hosting server555waps.cfa.harvard.edu/MIST/model_grids.html. We use the pre-packaged isochrone with v / vcrit = 0.4, and downloaded the basic version of the catalogue (25 columns). The MIST version number was 1.2, and the MESA revision number was 7503.
The main advantage of the MIST isochrones compared to older varieties is that they cover all phases of stellar evolution including remnant stars, which can be important for older stellar populations since late phase evolution and remnant stars are what provide the UV upturn component of galaxy SEDs. Missing these entirely (as many older isochrones do) can cause fitting issues for older stellar populations in highly quenched systems, and previous efforts to generate SSPs had to develop their own bespoke methods to create realistic upturns (Maraston & Thomas, 2000; Bruzual & Charlot, 2003). Notably some SSPs do not produce an upturn for older stellar populations, e.g. Maraston (2005) (as seen in Figure 14, discussed in detail later in this work).
@@ 3.1.2 PARSEC
PARSEC is a popular modern set of isochrones that really combine the efforts of two main groups: the Padova and Trieste isochrones that cover the majority of stellar evolution (e.g. the main sequence of stellar evolution) and later phases via the COLIBRI tracks (see Bressan et al., 2012; Marigo et al., 2017, for details). The latter covers the major asymptotic giant branch (AGB) phases that are extremely important for properly explaining the luminous output of infrared light in moderately old (few Gyr) old stars. Our compilation of the PARSEC isochrones was generated interactively via the main web server 666stev.oapd.inaf.it/cmd, and combines the v1.2S version of the Padova tracks with the S37 + S35 + PR16 combination of CALIBRI tracks (see the website for details on what these ingredients mean for the final isochrones). These isochrones cover ages in 0.1 dex bins between $10^{6.6}$$10^{10.1}$ years, and metallicity in 0.1 dex bins between $10^{-2.2}$$10^{0.5}$ $\text{Z}_{\odot}$. This means they generally cover a more limited range of ages and metallicity than MIST, and with coarser binning in age but finer in metallicity. In general for SED modelling, SSPs with finer and broader age sampling is advantageous.
The main limitation with PARSEC, and why they are not the recommended default set of isochrones, is that they lack proper remnant stellar evolution tracking. This is noted on the website, with a suggestion that this will be added longer term. As and when remnant stars are added to PARSEC, we will updated the standard isochrones available in ProGeny. We note that it is possible for users to adapt the provided isochrones and bolt on their own prescriptions for remnant star evolution (e.g. the Geneva isochrones, or the appropriate part of MIST). Counter to other SPLs, ProGeny avoids bespoke treatments for such evolutionary gaps. This means it relies more heavily on the evolutionary coverage of input isochrones, and lacks the explicit addition of features like planetary nebulae contributions that do exist in other libraries (Bruzual & Charlot, 2003).
@@ 3.1.3 BaSTI
The Bag of Stellar Tracks and Isochrones project (BaSTI) is an alternative isochrone library that monitors the evolutionary properties of stars up until their main sequence turn off (approximately). We downloaded pre-computed isochrones from the main INAF database777basti-iac.oa-abruzzo.inaf.it/isocs.html. The various files were converted into the appropriate format for ProGeny, with a particular note that the surface gravity properties of the stars had to be derived from the Stefan-Boltzmann Law (which means they are assumed to be non-rotating). These isochrones cover ages in 0.1 dex bins between $10^{7.3}$$10^{10.3}$ years, and metallicity in $\sim 0.2$ dex bins between $10^{-3.2}$$10^{0.45}$ $\text{Z}_{\odot}$. This means they generally cover a more limited range of ages and metallicity than MIST and PARSEC, and with coarser binning in age but finer in metallicity. The younger age limit is particular high (at roughly 20 Myr), making BaSTI a poor choice when computing properties of extreme star bursts (where sub 10 Myr sampling is critical).
@@ 3.1.4 Comparison
It is instructive to investigate some of the major differences between the MIST, PARSEC, and BaSTI isochrones included in the first released version of ProGeny. Figure 1 is a high-level comparison of the three families of isochrones at 100 Myr, 1 Gyr and 10 Gyr (the origin of the region polygons is made clear in Appendix A Figure 16). It is immediately apparent that the PARSEC and BaSTI isochrones are missing post-AGB stellar evolution including remnants. The main sequence region (labelled ‘MS’) is very similar between the three isochrone families, but the treatment of the RGB and TP-AGB region (bottom-left) does vary quite significantly, with PARSEC capturing the pulsations is relatively more detail. The more detailed tracking of the TP-AGB with PARSEC is not surprising, since it utilises the dedicated COLIBRI libraries for detailed modelling in this regime.
@@
The impact of variation in the TP-AGB region seems to be very modest in terms of the eventual impact on the final SSPs, but the lack of remnant stars in PARSEC and BaSTI has a very notable impact on the ultraviolet emission of older (few Gyr) stellar populations. If we apply our default interpolation of stellar spectra onto the isochrones we can compare the spectral flux prediction for a target Chabrier IMF. The result of this is shown in Figure 2, where there is a vastly different ultraviolet prediction due to the inclusion (MIST) or lack of inclusion (PARSEC, BaSTI) of hot remnant stars.
@@
As discussed later in Section 4, there is still a huge deal of uncertainty in the details of the ultraviolet upturn for old stellar populations and exactly what processes produce it (e.g. helium burning from Wolf-Rayet stars, post-AGB stars, horizontal branch stars, planetary nebulae, white dwarf remnants, and binary interactions; see Le Cras et al., 2016). There is also a scarcity of empirical data informing our knowledge of ultraviolet stellar spectra. The main conclusion to make here is that merely the inclusion or not of hot remnant stars (e.g. white dwarfs) alone will have a dramatic effect in the ultraviolet, but less impact at all on the optical to infrared regimes (i.e. any data long-wards of 2000 Å). Nebular emission can become significant in these wavelength regimes, but this will not be contributed to significantly by remnant stellar populations.
@@
Figure 2: Comparison of 10 Gyr ultraviolet upturn predictions for MIST, PARSEC, and BaSTI isochrones. The stellar spectra and interpolation used are otherwise identical, so the discrepancy in the ultraviolet is purely due to the inclusion (MIST) or lack of inclusion (PARSEC / BaSTI) of hot remnant stars. Note that the PARSEC and BaSTI spectra lie almost on top of each other in this plot.
3.2 Stellar Photosphere Spectra
@@
Equally as important as the isochrones, stellar spectra are the means by which we can convert our predictions of stellar evolution models to spectra that we can then use to model observed data, either directly as spectra, or passed through filters to create broadband photometry. Unlike isochrones, which are necessarily theoretical in nature (though informed by observations), stellar spectra can in principle be empirically derived using observations of many stars covering a large range of intrinsic properties, or generated theoretically. The main properties we usually wish to grid over are (in rough order of significance on the resulting spectrum): temperature; logarithmic surface gravity (referred to as ‘logG’ in the ProGeny package and this paper for compactness in Tables and Figures, where surface gravity has CGS units ($\text{cm}/\text{s}^{2}$) wherever mentioned); and metallicity (Z, always specified relative to solar in the ProGenypackage). More exotically, extensions are sometimes made to cover variations in stellar rotation and/or $\alpha$/Fe enhancement (or other abundance measures).
@@
In the perfect library of stellar spectra, there would exist a predicted or measured spectrum for every star for every combination of properties predicted by our stellar evolution isochrones, but this is unfeasible in practice — we would need millions of such spectra. The pragmatic solution is to map out a dense grid of these properties, with higher density where stars are most commonly found on the main sequence (since that will ultimately produce more accurate composite spectra). Rarer regions of temperate, surface gravity and metallicity can be covered more coarsely without biasing any resulting spectrum too much. Once we have established this dense grid, we can use the interpolation scheme (discussed above) to heuristically produce an approximate spectrum for any target combination of properties when moving along a given isochrone.
@@
Stellar population libraries exist that choose to place more emphasis on either empirical or theoretical / synthetic stellar spectra, but the main tradeoffs are accuracy (assuming we can estimate stellar properties, observations should always be closer to ‘reality’) and coverage. In the case of coverage, observed databases of stellar spectra tend to be much smaller (usually numbering hundreds), and only map out a limited range of parameter values. This makes interpolating or extrapolating these libraries onto target parameters quite complex, with spline interpolation and/or kriging schemes being necessary to achieve reasonable performance. In practice nearly all observationally focussed SPLs need to supplement the empirical stellar spectra with extreme/rare theoretical templates (e.g. for hot stars). The counter is also true, since some observed stellar spectra (e.g. TP_AGB stars) cannot be accurately reproduced theoretically.
@@
Table 3: Parameter ranges of the various stellar spectrum libraries initially includes with ProGeny. N is the total number of unique spectra.
Library Type Teff (K) logG logZ $\lambda$ (Å) N Reference
C3K (Conroy) Base 2,000 – 50,000 -1 – 5.5 -2.1 – 0.5 100 – $10^{8}$ 8,602 Conroy et al. (2018)
PHOENIX (Husser) Base 2,300 – 12,000 0 – 6 -4 – 1 500 – 55,000 7,559 Husser et al. (2013)
PHOENIX (Allard) Extend 2,000 – 70,000 0 – 5.5 -4 – 0.5 12.6 – 776,247 12,045 Allard et al. (2012)
MILES (Vazdekis) Alternative 2,000 – 50,000 -1 – 5.5 -1.4 – 0.2 3,525 – 7,500 3,915 Vazdekis et al. (2010)
BaSeL (WLBC) Alternative 2,000 – 50,000 -1 – 5.5 -2 – 0.5 91 – $1.6\times 10^{6}$ 4,649 Westera et al. (2002)
ATLAS9 (Castelli) Alternative 2,000 – 50,000 -1 – 5.5 -1 – 0.5 91 – $1.6\times 10^{6}$ 2,735 Castelli & Kurucz (2003)
AGB (Lacon) AGB 2,000 – 4,000 NA NA 5,100 – 24,900 14 Lançon & Mouhcine (2002)
TMAP (Werner) Hot 20,000 – 190,000 5 – 9 NA 3 – 400,000 124 Werner et al. (2003)
@@
Below we discuss the main decisions made regarding the fiducial stellar spectra used within ProGeny (see Table 3). It is always possible for users to supplement these with their own preferences and use the ProGeny software to generate new SSPs, but when we discuss the base ProGeny SSPs they have been constructed with a mixture of the below libraries. All stellar spectra provided as part of the fiducial ProGeny library encode the wavelength in angstroms and the luminosity in $\delta\lambda$ units such that the integral over all wavelengths of $f(\lambda).\delta\lambda=1$. This means all spectra can be scaled directly by the predicted stellar luminosity in an isochrone to achieve the correct output spectrum. Any additional stellar spectrum libraries added to ProGeny by users should obey the same requirements. The spectra are also re-binned to match the exact wavelength sampling of CB19, since via inspection this appeared to offer a good compromise of moderately high resolution (especially around important features), coverage and data volume. We note that when loading stellar spectra it is possible to target any desired wavelength grid, with flux conserving spectral resampling handled within ProGeny(written using efficient C++code).